The sum of two numbers is 9 and the sum of their squares is 261. Find the numbers
Answer by ewatrrr(24785)

x + y = 9 0r y = 9-x
....
(9-x)^2 + x^2 = 261
81 - 18x + x^2 + x^2 = 261
2x^2 - 18x - 180 = 0
x^2 - 9x - 90 = 0
(x+6)(x-15)= 0
x = 15 and y = -6
0r
x = -6 and y = 15
(15,-6) and (-6,15) solutions for this System
And..checking
225 + 36 = 261